# Populations of tidal and pulsating variables in eclipsing binaries

### v1

@@ 5 Conclusion
The positions of the hybrid and the p-mode pulsators in the HRD are as we expect for these pulsators, peaking strongly around the Sct region, with the hybrid pulsators having somewhat lower masses than the pure p-mode pulsators. The far more numerous g-mode pulsators peak in the Dor region and also populate the SPB region. However, there is no sign of a well-defined break between the two regions in the HRD. Instead, the distribution of g-mode pulsators is continuous towards higher masses. The g-mode pulsators are biased towards shorter orbital periods and towards higher values of our orbital-harmonic-based tidal morphology parameter, which is indicative of the strength of tidal variability in the light curve. However, this effect appears to be small. We conclude that tidal effects are unlikely to be the primary cause of the departures from the classical regions of instability found in this and previous works dealing with large populations of pulsators (Gaia Collaboration et al., 2023a; Hey & Aerts, 2024; Mombarg et al., 2024a), leaving the physical limitations of mode excitation theories and sample contamination by rotational variables as leading explanations.
@@
Eclipsing binaries featuring hybrid pulsations more closely match the g-mode distribution when it comes to orbital properties – and correlation with tidal variability – than the p-mode distribution, while closely matching the p-mode distribution in terms of surface properties such as log and log. The correlation with the p-mode properties is intuitive, given that the p-mode distribution is dominated by Sct pulsators and contains only a few (presumed) Cep pulsators. The preference for short-period, circular binaries for both pure g-mode pulsators and hybrid pulsators is consistent with more rapid spin-orbit coupling and circularisation in the presence of g-mode pulsations, presumably driven by increased tidal efficiency and angular momentum transport (Aerts et al., 2019).
@@ Acknowledgements.
The authors wish to acknowledge useful discussions with Dario Fritzewski, Haotian Wang, and Vincent Vanlaer. The authors also wish to thank the anonymous referee for their constructive report. The research leading to these results has received financial support from the Flemish Government under the long-term structural Methusalem funding program by means of the project SOUL: Stellar evolution in full glory, grant METH/24/012, and from the European Research Council (ERC) under the Horizon Europe programme (Synergy Grant agreement No. 101071505: 4D-STAR). While partially funded by the European Union, views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The authors also acknowledge the Belgian Federal Science Policy Office (BELSPO) for their financial support in the framework of the PRODEX Programme of the European Space Agency (ESA), facilitating the exploitation of the Gaia data.References
- Aerts (2021) Aerts, C. 2021, Reviews of Modern Physics, 93, 015001
- Aerts et al. (2019) Aerts, C., Mathis, S., & Rogers, T. M. 2019, ARA&A, 57, 35
- Aerts et al. (2018) Aerts, C., Molenberghs, G., Michielsen, M., et al. 2018, ApJS, 237, 15
- Aerts & Tkachenko (2024) Aerts, C. & Tkachenko, A. 2024, A&A, 692, R1
- Aerts et al. (2025) Aerts, C., Van Reeth, T., Mombarg, J. S. G., & Hey, D. 2025, A&A, 695, A214
- Audenaert et al. (2021) Audenaert, J., Kuszlewicz, J. S., Handberg, R., et al. 2021, AJ, 162, 209
- Audenaert & Tkachenko (2022) Audenaert, J. & Tkachenko, A. 2022, A&A, 666, A76
- Balona (2023) Balona, L. A. 2023, arXiv e-prints, arXiv:2310.09805
- Borucki et al. (2010) Borucki, W. J., Koch, D., Basri, G., et al. 2010, Science, 327, 977
- Bouabid et al. (2013) Bouabid, M. P., Dupret, M. A., Salmon, S., et al. 2013, MNRAS, 429, 2500
- Bowman et al. (2016) Bowman, D. M., Kurtz, D. W., Breger, M., Murphy, S. J., & Holdsworth, D. L. 2016, MNRAS, 460, 1970
- Castelli & Kurucz (2003) Castelli, F. & Kurucz, R. L. 2003, in IAU Symposium, Vol. 210, Modelling of Stellar Atmospheres, ed. N. Piskunov, W. W. Weiss, & D. F. Gray, A20
- Choi et al. (2016) Choi, J., Dotter, A., Conroy, C., et al. 2016, ApJ, 823, 102
- Dupret et al. (2005) Dupret, M. A., Grigahcène, A., Garrido, R., Gabriel, M., & Scuflaire, R. 2005, A&A, 435, 927
- Dziembowski et al. (1993) Dziembowski, W. A., Moskalik, P., & Pamyatnykh, A. A. 1993, MNRAS, 265, 588
- Dziembowski & Pamiatnykh (1993) Dziembowski, W. A. & Pamiatnykh, A. A. 1993, MNRAS, 262, 204
- Freytag et al. (1996) Freytag, B., Ludwig, H. G., & Steffen, M. 1996, A&A, 313, 497
- Fritzewski et al. (2024) Fritzewski, D. J., Van Reeth, T., Aerts, C., et al. 2024, A&A, 681, A13
- Fuller (2017) Fuller, J. 2017, MNRAS, 472, 1538
- Fuller & Lai (2012) Fuller, J. & Lai, D. 2012, MNRAS, 420, 3126
- Gaia Collaboration et al. (2023a) Gaia Collaboration, De Ridder, J., Ripepi, V., et al. 2023a, A&A, 674, A36
- Gaia Collaboration et al. (2023b) Gaia Collaboration, Vallenari, A., Brown, A. G. A., et al. 2023b, A&A, 674, A1
- Gautschy & Saio (1993) Gautschy, A. & Saio, H. 1993, MNRAS, 262, 213
- Guo (2021) Guo, Z. 2021, Frontiers in Astronomy and Space Sciences, 8, 67
- Guo et al. (2019) Guo, Z., Fuller, J., Shporer, A., et al. 2019, ApJ, 885, 46
- Guzik et al. (2000) Guzik, J. A., Kaye, A. B., Bradley, P. A., Cox, A. N., & Neuforge, C. 2000, ApJ, 542, L57
- Hambleton et al. (2013) Hambleton, K. M., Kurtz, D. W., Prša, A., et al. 2013, MNRAS, 434, 925
- Herwig (2000) Herwig, F. 2000, A&A, 360, 952
- Hey & Aerts (2024) Hey, D. & Aerts, C. 2024, A&A, 688, A93
- Hodges (1958) Hodges, J. L. 1958, Arkiv for Matematik, 3, 469
- Hon et al. (2024) Hon, M., Li, Y., & Ong, J. 2024, ApJ, 973, 154
- Huang et al. (2020a) Huang, C. X., Vanderburg, A., Pál, A., et al. 2020a, Research Notes of the American Astronomical Society, 4, 204
- Huang et al. (2020b) Huang, C. X., Vanderburg, A., Pál, A., et al. 2020b, Research Notes of the American Astronomical Society, 4, 206
- Hubeny & Lanz (2017) Hubeny, I. & Lanz, T. 2017, arXiv e-prints, arXiv:1706.01859
- IJspeert et al. (2024a) IJspeert, L. W., Tkachenko, A., Johnston, C., & Aerts, C. 2024a, A&A, 691, A242
- IJspeert et al. (2024b) IJspeert, L. W., Tkachenko, A., Johnston, C., et al. 2024b, A&A, 685, A62
- Kemp et al. (2025) Kemp, A., Fritzewski, D. J., Van Reeth, T., et al. 2025, A&A, 693, A184
- Kemp et al. (2024) Kemp, A., Tkachenko, A., Torres, G., et al. 2024, A&A, 689, A164
- Kirk et al. (2016) Kirk, B., Conroy, K., Prša, A., et al. 2016, AJ, 151, 68
- Kunimoto et al. (2021) Kunimoto, M., Huang, C., Tey, E., et al. 2021, Research Notes of the American Astronomical Society, 5, 234
- Kurtz et al. (2014) Kurtz, D. W., Saio, H., Takata, M., et al. 2014, MNRAS, 444, 102
- Lanz & Hubeny (2003) Lanz, T. & Hubeny, I. 2003, ApJS, 146, 417
- Lanz & Hubeny (2007) Lanz, T. & Hubeny, I. 2007, ApJS, 169, 83
- Massey Jr (1951) Massey Jr, F. J. 1951, Journal of the American statistical Association, 46, 68
- Michielsen (2024) Michielsen, M. 2024, Journal of Open Source Software, 9, 5884
- Michielsen et al. (2021) Michielsen, M., Aerts, C., & Bowman, D. M. 2021, A&A, 650, A175
- Michielsen et al. (2023) Michielsen, M., Van Reeth, T., Tkachenko, A., & Aerts, C. 2023, A&A, 679, A6
- Miglio et al. (2007) Miglio, A., Montalbán, J., & Dupret, M.-A. 2007, MNRAS, 375, L21
- Mombarg et al. (2024a) Mombarg, J. S. G., Aerts, C., & Molenberghs, G. 2024a, A&A, 685, A21
- Mombarg et al. (2024b) Mombarg, J. S. G., Aerts, C., Van Reeth, T., & Hey, D. 2024b, A&A, 691, A131
- Mombarg et al. (2020) Mombarg, J. S. G., Dotter, A., Van Reeth, T., et al. 2020, ApJ, 895, 51
- Murphy et al. (2024) Murphy, S. J., Bedding, T. R., Gautam, A., Kerr, R. P., & Mani, P. 2024, MNRAS, 534, 3022
- Murphy et al. (2019) Murphy, S. J., Hey, D., Van Reeth, T., & Bedding, T. R. 2019, MNRAS, 485, 2380
- Ogilvie (2014) Ogilvie, G. I. 2014, ARA&A, 52, 171
- Pedersen et al. (2021) Pedersen, M. G., Aerts, C., Pápics, P. I., et al. 2021, Nature Astronomy, 5, 715
- Prša et al. (2011) Prša, A., Batalha, N., Slawson, R. W., et al. 2011, AJ, 141, 83
- Prša et al. (2022) Prša, A., Kochoska, A., Conroy, K. E., et al. 2022, ApJS, 258, 16
- Rauer et al. (2025) Rauer, H., Aerts, C., Cabrera, J., et al. 2025, Experimental Astronomy, 59, 26
- Ricker et al. (2015) Ricker, G. R., Winn, J. N., Vanderspek, R., et al. 2015, Journal of Astronomical Telescopes, Instruments, and Systems, 1, 014003
- Rogers et al. (2013) Rogers, T. M., Lin, D. N. C., McElwaine, J. N., & Lau, H. H. B. 2013, ApJ, 772, 21
- Schmid & Aerts (2016) Schmid, V. S. & Aerts, C. 2016, A&A, 592, A116
- Sekaran et al. (2021) Sekaran, S., Tkachenko, A., Johnston, C., & Aerts, C. 2021, A&A, 648, A91
- Skrutskie et al. (2006) Skrutskie, M. F., Cutri, R. M., Stiening, R., et al. 2006, AJ, 131, 1163
- Slawson & Doyle (2011) Slawson, R. W. & Doyle, L. R. 2011, in American Astronomical Society Meeting Abstracts, Vol. 217, American Astronomical Society Meeting Abstracts #217, 140.12
- Stankov & Handler (2005) Stankov, A. & Handler, G. 2005, ApJS, 158, 193
- Uytterhoeven et al. (2011) Uytterhoeven, K., Moya, A., Grigahcène, A., et al. 2011, A&A, 534, A125
- Waelkens (1991) Waelkens, C. 1991, A&A, 246, 453
- Winkler et al. (2019) Winkler, C., Worrall, D., Hoogeboom, E., & Welling, M. 2019, arXiv e-prints, arXiv:1912.00042
Appendix A Additional tables and figures

### v2

@@ 5 Conclusion
The positions of the hybrid and the p-mode pulsators in the HRD are as we expect for these pulsators, peaking strongly around the Sct region, with the hybrid pulsators having somewhat lower masses than the pure p-mode pulsators. The far more numerous g-mode pulsators peak in the Dor region and also populate the SPB region. However, there is no sign of a well-defined break between the two regions in the HRD. Instead, the distribution of g-mode pulsators is continuous towards higher masses. The g-mode pulsators are biased towards shorter orbital periods and towards higher values of our orbital-harmonic-based tidal morphology parameter, which is indicative of the strength of tidal variability in the light curve. However, this effect appears to be small. We conclude that tidal effects are unlikely to be the primary cause of the departures from the classical regions of instability found in this and previous works dealing with large populations of pulsators (gaia2023pulsations; hey2024; mombarg2024), leaving the physical limitations of mode excitation theories and sample contamination by rotational variables as leading explanations.
@@
Eclipsing binaries featuring hybrid pulsations more closely match the g-mode distribution when it comes to orbital properties – and correlation with tidal variability – than the p-mode distribution, while closely matching the p-mode distribution in terms of surface properties such as log and log. The correlation with the p-mode properties is intuitive, given that the p-mode distribution is dominated by Sct pulsators and contains only a few (presumed) Cep pulsators. The preference for short-period, circular binaries for both pure g-mode pulsators and hybrid pulsators is consistent with more rapid spin-orbit coupling and circularisation in the presence of g-mode pulsations, presumably driven by increased tidal efficiency and angular momentum transport (aerts2019).
@@ Acknowledgements.
The authors wish to acknowledge useful discussions with Dario Fritzewski, Haotian Wang, and Vincent Vanlaer. The authors also wish to thank the anonymous referee for their constructive report. The research leading to these results has received financial support from the Flemish Government under the long-term structural Methusalem funding program by means of the project SOUL: Stellar evolution in full glory, grant METH/24/012, and from the European Research Council (ERC) under the Horizon Europe programme (Synergy Grant agreement No. 101071505: 4D-STAR). While partially funded by the European Union, views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The authors also acknowledge the Belgian Federal Science Policy Office (BELSPO) for their financial support in the framework of the PRODEX Programme of the European Space Agency (ESA), facilitating the exploitation of the Gaia data.Appendix A Additional tables and figures